40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  ATF eLog  Not logged in ELOG logo
Entry  Tue Sep 14 04:14:25 2010, Dmass, Misc, Doubling, Doubling results summary 
    Reply  Wed Sep 15 17:04:36 2010, Dmass, Misc, Doubling, DON'T REMOVE OPTICS YET 
       Reply  Wed Sep 15 22:59:37 2010, Frank, Misc, Doubling, DON'T REMOVE OPTICS YET 
    Reply  Tue Sep 21 01:52:05 2010, Dmass, Misc, Doubling, Doubling results summary AllanVarSep2.pdf
       Reply  Thu Oct 6 23:44:09 2011, Dmass, Misc, Doubling, Finishing SHG paper AvarWhitePhase.pdfAvarWhiteFreqNoise.pdfAvarPhase1onfhalf.pdf
          Reply  Mon Oct 10 11:55:59 2011, Dmass, Misc, Doubling, Finishing SHG paper Avar2PhaseSchnatz.pdf
    Reply  Wed Sep 22 17:16:09 2010, Dmass, Misc, Doubling, Some notes on previous bounds 
       Reply  Mon Jan 24 17:17:59 2011, Dmass, Misc, Doubling, Some notes on previous bounds 
Message ID: 1053     Entry time: Tue Sep 14 04:14:25 2010     Reply to this: 1054   1074   1079
Author: Dmass 
Type: Misc 
Category: Doubling 
Subject: Doubling results summary 

This post will detail the end results of the Mach Zehnder doubling noise experiment.

  • I made a full noise budget with the best measurement of Mach Zehnder phase noise that I'm going to get.
  • I broke off the leads to one of the oven's temperature sensors.
  • At some point after this, one of the Newport 3040 Temperature Controllers failed.
  • I added some NTC 833 thermistors to each oven, with beads of epoxy on top of each one
    • I make a 4 lead measurement using these with one AD587using 100 kOhm resistors in series - I pull ~ 130 uA from it
    • I use an SR560 to AC couple the signal from the NTC 833's with a gain of 500 and a bandpass from 0.03 Hz to 10 Hz (one pole one zero)
  • I took the control signal from the working 3040, and used it to drive the other ovens TEC
  • I detuned the oven with the working 3040 so that both ovens produced ~ the same power in green.
    • This means that neither of the ovens are at the peak of their phase matching curves, which has unknown effects.
  • I put a line on the laser intensity at 85 Hz to see what the intensity noise coupling is
  • I turned the HEPA on and off using its lowest setting, getting these results
    • The low frequency noise (below 1 Hz) seems to go down with the lowest HEPA setting (this doesn't seem impossible)
    • The noise above 1 Hz significantly increases
    • For the subtraction, the intensity noise line shows up at about 23 dB above the background, I think we are intensity noise limited but need to double check the line height (easy to do).
    • The high frequency limit is much higher than the results in the "full noise budget", so even if it is intensity limit, it could be the oven detuning which changes this coupling drastically. I can't say that in the good measurement we are limited at high frequency by intensity noise.
    • I believe we are limited by air currents below 10 Hz
    • I think that the noise going up below 1 Hz when we turn the HEPA to "OFF" is not inconsistent with being limited by air currents. It's just like asking "if you took a fluid, could you decrease the temperature variation at one point by stirring it". I think the answer is maybe. I also think that thinking about it too hard will be a waste of time, and that the result is "temperature fluctuations from air currents are the limiting noise coupling, and better ovens / a better box (vacuum) would fix this.

Update - I have uploaded an omnigraffle diagram onto the svn. /trunk/docs/DoublingNoise/MZFullDiagram.graffle. I have tried to make it semi comprehensive in terms of gains, settings, and part numbers.

ELOG V3.1.3-